OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet?

OpenStudy (anonymous):

width = x length = twice of width = 2x height = 3 less than width = x-3 now use the fact that volume = length*width*height what do you get ????

OpenStudy (anonymous):

4x - 3? I'm not sure on how to do this i'm horrible with math...

OpenStudy (anonymous):

vol = (x)*(2x)*(x-3) vol = (2x^2)*(x-3) vol = 2x^3 - 6x^2

OpenStudy (anonymous):

Okay got it thank you...

OpenStudy (anonymous):

u r welcome...☺